Check my work Check My Work button is now enabledItem 7Item 7 6.67 points Item Skipped Solo Corp. is evaluating a project with the following cash flows: Year Cash Flow 0 –$ 29,800 1 12,000 2 14,700 3 16,600 4 13,700 5 – 10,200 The company uses a discount rate of 13 percent and a reinvestment rate of 6 percent on all of its projects. a. Calculate the MIRR of the project using the discounting approach. (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.) b. Calculate the MIRR of the project using the reinvestment approach. (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.) c. Calculate the MIRR of the project using the combination approach. (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

a. Discounting approach MIRR %
b. Reinvestment approach MIRR %
c. Combination approach MIRR %
Answer #1

Discount Rate = 13%
Reinvestment Rate = 6%

Answer a.

Present Value of Cash Outflow = $29,800 + $10,200/1.13^5
Present Value of Cash Outflow = $35,336.15135

Let MIRR be i%

0 = -$35,336.15135 + $12,000/(1+i) + $14,700/(1+i)^2 + $16,600/(1+i)^3 + $13,700/(1+i)^4

Using financial calculator, i = 21.56%

MIRR using discounting method is 21.56%

Answer b.

Future Value of Cash Flows = $12,000*1.06^4 + $14,700*1.06^3 + $16,600*1.06^2 + $13,700*1.06 - $10,200
Future Value of Cash Flows = $55,631.41872

Let MIRR be i%

0 = -$29,800 + $55,631.41872/(1+i)^5
(1+i)^5 = 1.866826
1+i = 1.1330
i = 0.1330 or 13.30%

MIRR using reinvestment method is 13.30%

Answer c.

Present Value of Cash Outflow = $29,800 + $10,200/1.13^5
Present Value of Cash Outflow = $35,336.15135

Future Value of Cash Inflows = $12,000*1.06^4 + $14,700*1.06^3 + $16,600*1.06^2 + $13,700*1.06
Future Value of Cash Inflows = $65,831.41872

MIRR = (Future Value of Cash Inflows / Present Value of Cash Outflow)^(1/Period) - 1
MIRR = ($65,831.41872 / $35,336.15135)^(1/5) - 1
MIRR = 1.8630^(1/5) - 1
MIRR = 1.1325 - 1
MIRR = 0.1325 or 13.25%

MIRR using combination method is 13.25%